# ERDŐS–PÓSA PROPERTY OF CYCLES THAT ARE FAR APART

### v2

@@ Acknowledgements
In this appendix, we present a complete proof of Theorem 3 and show that, for any fixed $c$, the bounding function $\ell^{\star}(k,c)$ is polynomial in $k$. All of the ideas from this proof come from Gyárfás and Lehel [4], who proved it when the host forest $F$ is a path and sketched changes needed for the general case.
@@
Let $F$ be a forest. For each component of $F$, fix an arbitrary vertex of $F$. This makes $F$ a rooted forest. For each component $C$ in an arbitrary subgraph $H$ of $F$, there is a unique vertex $v_{C}$ that is closest to a root of $F$. We call $v_{C}$ the root of $C$ in $F$. We treat $H$ as a rooted forest, where each component $C$ of $H$ is rooted at $v_{C}$. For two vertices $u$ and $v$ in $F$, if $v$ is contained in the unique path from $u$ to a root of $F$, then we say that $u$ is a descendant of $v$ in $F$, and when additionally $u\neq v$ then $u$ is a strict descendant of $v$ in $F$. If $u$ is a strict descendant of $v$ and $uv\in E(F)$ then $u$ is a child of $v$. For a vertex $v$ in $F$ the subtree, $\leavevmode\color[rgb]{0.76,0.13,0.28}F(v)$, of $v$ is the subgraph of $F$ induced by all the descendants of $v$ in $F$ that is rooted at $v$. Let $\sigma=(v_{1},\ldots,v_{n})$ be an ordering of the vertices of $F$. We say that $u$ precedes $v$ in $\sigma$ if $u=v_{i}$, $v=v_{j}$ and $i\leq j$. We say that $\sigma$ is a topological ordering of $F$ if for each $v\in V(F)$, every vertex of $F(v)$ precedes $v$ in $\sigma$.
@@
For an indexed family $A:=(A_{i})_{i\in I}$ with index set $I$, we use $\leavevmode\color[rgb]{0.76,0.13,0.28}\pi_{i}(A)$ to denote the element $A_{i}$ of $A$ indexed by $i$. Let $I$ be a set and let $F_{i}$ be a forest for each $i\in I$ [^4] In our application, the forests in $(F_{i})_{i\in I}$ are pairwise vertex-disjoint, but this is not required in the following results. A sequence $(A_{i})_{i\in I}$ is a $(F_{i})_{i\in I}$-tuple if, for each $i\in I$, $A_{i}$ is a null graph or $A_{i}$ is a connected subgraph of $F_{i}$. An $(F_{i})_{i\in I}$-tuple $(A_{i})_{i\in I}$ is trivial if $A_{i}$ is a null graph for each $i\in I$ and is non-trivial otherwise. Let $A=(A_{i})_{i\in I}$ and $B=(B_{i})_{i\in I}$ be two $(F_{i})_{i\in I}$-tuples. We say that $A$ and $B$ are independent if $V(A_{i})\cap V(B_{i})=\emptyset$ for each $i\in I$. A collection $\mathcal{A}$ of $(F_{i})_{i\in I}$-tuples is independent if the elements of $\mathcal{A}$ are pairwise independent.
@@
We make use of the following corollary, which is obtained immediately from Lemma 14, below, by taking $m=k$ and $x_{1}=x_{2}=\cdots=x_{k}=1$.
@@
Let $c,k$ be positive integers. Let $F_{i}$ be a forest for each $i\in[c]$. For each $j\in[k]$, let $\mathcal{A}_{j}$ be a collection of pairwise-independent $(F_{i})_{i\in[c]}$-tuples with $|\mathcal{A}_{j}|\geq k^{c}$. Then, for each $j\in[k]$ there exists $B_{j}\in\mathcal{A}_{j}$ such that $B_{1},\ldots,B_{k}$ are pairwise independent.
To establish 13 we prove the following more general result, which is amenable to a proof by induction.
@@
Let $c,m,k$ be positive integers. Let $F_{i}$ be a forest for each $i\in[c]$. For each $j\in[m]$, let $\mathcal{A}_{j}$ be a collection of pairwise-independent $(F_{i})_{i\in[c]}$-tuples with $|\mathcal{A}_{j}|\geq m^{c-1}k$. Let $x_{1},\ldots,x_{m}$ be nonnegative integers with $x_{1}+\cdots+x_{m}=k$. Then, for each $j\in[m]$ there exist $A_{j}^{1},\ldots,A_{j}^{x_{j}}\in\mathcal{A}_{j}$ such that
@@
We call a tuple $(c,m,k,(F_{i})_{i\in[c]},(\mathcal{A}_{j})_{j\in[m]},(x_{j})_{j\in[m]})$ that satisfies the assumptions of the lemma a $(c,m,k)$-instance and we say that the sequence of sets in Equation 10 satisfies this $(c,m,k)$-instance.
@@
Without loss of generality, we may assume that $x_{j}>0$ for each $j\in[m]$. To see why this is so, let $J:=\{j\in[m]:x_{j}>0\}$ and $m^{\prime}:=|J|$. Since $m^{\prime}\leq m$, $\mathcal{I}^{\prime}:=(c,m^{\prime},k,(F_{i})_{i\in[c]},(\mathcal{A}_{j})_{j\in J},(x_{j})_{j\in J})$ is a $(c,m^{\prime},k)$-instance with $x_{j}>0$ for each $j\in J$, and any sequence $(A_{j}^{\ell})_{j\in I,\,\ell\in[x_{j}]}$ that satisfies this instance also satisfies Equation 10. From this point on we assume that $x_{j}>0$ for each $j\in[m]$.
@@
Without loss of generality, we assume that for all $j\in[k]$, $i\in[c]$, $A\in\mathcal{A}_{j}$, the tree $\pi_{i}(A)$ is non-null. Indeed, otherwise for all $j$, $i$, $A$ violating this rule, we introduce a new vertex in $F_{i}$ and make $\pi_{i}(A)$ be the graph that contains only this new vertex. Since this new vertex appears only in $\pi_{i}(A)$, the set $\mathcal{A}_{j}$ is still independent. Because $\mathcal{A}_{j}$ is independent for each $j\in[m]$, this assumption ensures that $\pi_{i}(A)\neq\pi_{i}(A^{\prime})$ for each $i\in[c]$, $j\in[m]$ and distinct $A,A^{\prime}\in\mathcal{A}_{j}$. This ensures that, for every non-empty $I\subseteq[c]$, every $j\in[m]$, and every $\tilde{\mathcal{A}}_{j}\subseteq\mathcal{A}_{j}$, we have |$|\{(\pi_{i}(A))_{i\in I}:A\in\tilde{\mathcal{A}}_{j}\}|=|\tilde{\mathcal{A}}_{j}|$. In the following we will use this fact repeatedly without explicitly referring to it.
@@
This assumption has another consequence, which we use to simplify the description of sequence in Equation 10. Let $I$ be a non-empty subset of $[m]$, let $\mathcal{B}_{j}\subseteq\mathcal{A}_{j}$ for each $j\in[m]$, and let $\mathcal{I}_{\mathcal{B}}:=(|I|,m,k^{\prime},(F_{i})_{i\in I},(\mathcal{B}_{j})_{j\in[m]},(x^{\prime}_{j})_{j\in[m]})$ be a $(|I|,m^{\prime},k^{\prime})$-instance. Then, for each $B\in\bigcup_{j\in[m]}\mathcal{B}_{j}$, the $(F_{i})_{i\in I}$-tuple $(\pi_{i}(B))_{i\in I}$ is non-trivial. Any two non-trivial independent $(F_{i})_{i\in I}$-tuples are necessarily distinct. Therefore, any sequence that satisfies $\mathcal{I}_{\mathcal{B}}$ is a sequence of $k^{\prime}$ distinct $(F_{i})_{i\in I}$-tuples. Treating this sequence as a set does not change its size. Conversely, if there exist sets $\mathcal{B}_{1}^{\prime},\ldots,\mathcal{B}_{m}^{\prime}$ with $\mathcal{B}_{j}^{\prime}\subseteq\mathcal{B}_{j}$ and $|\mathcal{B}_{j}^{\prime}|=x_{j}^{\prime}$ for each $j\in[m]$ and such that $\bigcup_{j\in[m]}\mathcal{B}_{j}^{\prime}$ is independent and has size $k^{\prime}$ then there exists a sequence that satisfies $\mathcal{I}_{\mathcal{B}}$. In this case, we say that $\mathcal{B}_{1}^{\prime},\ldots,\mathcal{B}_{m}^{\prime}$ satisfies $\mathcal{I}_{\mathcal{B}}$.
@@
With these two assumptions justified, we now proceed with the proof. Let $\mathcal{I}:=(c,m,k,(F_{i})_{i\in[c]},(\mathcal{A}_{j})_{j\in[m]},(x_{j})_{j\in[m]})$ be a $(c,m,k)$-instance satisfying the assumptions in the previous paragraphs. Our goal is to show the existence of $\mathcal{A}^{\prime}_{1},\ldots,\mathcal{A}^{\prime}_{m}$ that satisfies $\mathcal{I}$. We proceed by induction on $(c,k)$ in lexicograph order.
@@
Define $y_{j}:=x_{j}$ for each $j\in[m]\setminus\{\ell\}$ and $y_{\ell}:=x_{\ell}-1$. Now, (11) implies that $\mathcal{I}_{\mathcal{B}}:=(1,m,k-1,(F_{i})_{i\in[c]},(\mathcal{B}_{j})_{j\in[m]},(y_{j})_{j\in[m]})$ is an $(1,m,k-1)$-instance. By the inductive hypothesis, there exists $\mathcal{B}_{1}^{\prime},\ldots,\mathcal{B}_{m}^{\prime}$ that satisfy $\mathcal{I}_{\mathcal{B}}$. For each $j\in[m]\setminus\{\ell\}$, let $\mathcal{A}^{\prime}_{j}:=\mathcal{B}^{\prime}_{j}$ and let $\mathcal{A}^{\prime}_{\ell}:=\{\check{A}\}\cup\mathcal{B}^{\prime}_{\ell}$. Then the sequence $\mathcal{A}^{\prime}_{1},\ldots,\mathcal{A}^{\prime}_{m}$ satisfies $\mathcal{I}$. This completes the proof of the case $c=1$.
@@
is a $(1,m,m^{c-1}k)$-instance. Applying induction on $\mathcal{I}_{\mathcal{B}}$ we obtain $\mathcal{B}_{1}^{\prime},\ldots,\mathcal{B}_{m}^{\prime}$ that satisfies $\mathcal{I}_{\mathcal{B}}$.
For each $j\in[m]$, let $\mathcal{C}_{j}=\{(\pi_{2}(A),\ldots,\pi_{c}(A)):A\in\mathcal{A}_{j},\pi_{1}(A)\in\mathcal{B}^{\prime}_{j}\}$. Observe that $|\mathcal{C}_{j}|=|\mathcal{B}^{\prime}_{j}|=y_{j}=m^{c-2}k$. Therefore, the tuple $\mathcal{I}_{\mathcal{C}}:=(c-1,m,k,(F_{i})_{i\in\{2,\ldots,c\}},(\mathcal{C}_{j})_{j\in[m]},(x_{j})_{j\in[m]})$ is a $(c-1,k)$-instance. Applying induction yields $\mathcal{C}_{1}^{\prime},\ldots,\mathcal{C}_{m}^{\prime}$ that satisfies $\mathcal{I}_{\mathcal{C}}$.
@@
There exists a function $\ell:\mathbb{N}^{2}\to\mathbb{N}$ such that, for all $k,c\in\mathbb{N}$ with $k,c\geq 1$, for every $(F_{i})_{i\in[c]}$ where $F_{i}$ is a forest for each $i\in[c]$, for every collection $\mathcal{A}$ of non-trivial $(F_{i})_{i\in[c]}$-tuples either
@@
We may assume that $\pi_{i}(A)$ contains at least one vertex, for each $A\in\mathcal{A}$ and each $i\in[c]$. Indeed, otherwise we introduce a new vertex in $F_{i}$ and make $\pi_{i}(A)$ be the graph that contains only this new vertex. Since this vertex appears only in $\pi_{i}(A)$, this does not affect outcome a. If this new vertex appears in the set $X_{i}$ promised by outcome b, then we can remove it from $X_{i}$ and replace it with any other vertex in $\bigcup_{j\in[c]\setminus\{i\}}V(\pi_{j}(A))$. (The fact that $A$ is non-trivial for each $A\in\mathcal{A}$ ensures that $\bigcup_{j\in[c]\setminus\{i\}}V(\pi_{j}(A))$ contains at least one vertex.) This assumption ensures that for any non-empty $I\subseteq[c]$, and any $A\in\mathcal{A}$, $(\pi_{i}(A))_{i\in I}$ is non-trivial.
@@
Finally, we may assume that for each $i\in[c]$, $F_{i}$ is a binary tree, [^5] A binary tree is a rooted tree in which each vertex has at most two children. The root of a binary tree is not considered to be a leaf, even if it has no children. by the following argument. For each $i\in[c]$, and each vertex $v$ in $F_{i}$ with children $u_{1},\ldots,u_{\beta}$, define $T_{i}(v)$ to be a binary tree rooted at $v$ and whose leaves are $u_{1},\ldots,u_{\beta}$, and let $T^{-}_{i}(v):=T_{i}(v)-\{u_{1},\ldots,u_{\beta}\}$. Now let $F_{i}^{\prime}:=\bigcup_{v\in V(F_{i})}T_{i}(v)$ and note that $F_{i}^{\prime}$ is a binary tree for each $i\in[c]$. For each $A\in\mathcal{A}$, define $A^{\prime}:=(A^{\prime}_{1},\ldots,A^{\prime}_{c})$ where $A^{\prime}_{i}:=F_{i}^{\prime}[\bigcup_{v\in V(A)}V(T^{-}_{i}(v))]$ for each $i\in[c]$ and define $\mathcal{A}^{\prime}:=\{A^{\prime}:A\in\mathcal{A}\}$. Then $\mathcal{A}^{\prime}$ is a collection of non-trivial $(F_{i}^{\prime})_{i\in[c]}$-tuples. Observe that for all $A,B\in\mathcal{A}$, (1) $A$and $B$ are independent if and only if $A^{\prime}$ and $B^{\prime}$ are independent; (2) if $A_{i}^{\prime}$ contains a vertex in $T_{i}^{-}(v)$ for some $i\in[c]$ and $v\in V(F_{i})$, then $A_{i}$ contains $v$. Now apply the lemma to $(F^{\prime}_{i})_{i\in[c]}$ and $\mathcal{A}^{\prime}$. If this results in an independent set $I^{\prime}\subseteq\mathcal{A}^{\prime}$ of size $k$, then Item 1 ensures that $I:=\{A\in\mathcal{A}:A^{\prime}\in I^{\prime}\}$ satisfies a. If this results in sets $X_{1}^{\prime},\ldots,X_{c}^{\prime}$, then define $X_{i}:=\{v\in V(F_{i}):V(T_{i}^{-}(v))\cap X_{i}^{\prime}\neq\emptyset\}$ for each $i\in[c]$. Then Item 2 ensures that the sets $X_{1},\ldots,X_{c}$ satisfy b.
We now proceed by induction on $c$. In order to avoid presenting an argument for $c=1$ that is a special case of the argument for $c>1$, we treat $c=0$ as the base case of the induction. In this case, the conclusion of the lemma is trivial since the empty sequence satisfies the conditions of outcome b.
Now suppose that $c\geq 1$. Let $n=|V(F_{1})|$, let $\sigma$ be a topological ordering of the vertices of $F_{1}$ and, for $v,w\in V(F_{1})$ use the notation $v<w$ if and only if $v\neq w$ and $v$ precedes $w$ in $\sigma$. For convenience, we define an artificial vertex $v_{0}$ such that $v_{0}<v$ for all $v\in V(F_{1})$. Consider the following iterative process, that constructs a sequence $v_{1}<\cdots<v_{m}$ of vertices in $V(F_{1})$ and a sequence $\mathcal{D}_{1},\ldots,\mathcal{D}_{m}$ of $(F_{i})_{i\in\{2,\ldots,c\}}$-tuples. Suppose that $v_{1},\ldots,v_{z-1}$ and $\mathcal{D}_{1},\ldots,\mathcal{D}_{z-1}$ have already been defined, for some $z\geq 1$ and define

### v3

@@ 5 Proof of Theorem˜3
In this section, we present a complete proof of Theorem˜3 and show that, for any fixed $c$, the bounding function $\ell^{\star}(k,c)$ is polynomial in $k$. All of the ideas from this proof come from Gyárfás and Lehel [4], who proved it when the host forest $F$ is a path and sketched changes needed for the general case.
@@
Let $F$ be a forest. For each component of $F$, fix an arbitrary vertex of $F$. This makes $F$ a rooted forest. For each component $C$ in an arbitrary subgraph $H$ of $F$, there is a unique vertex $v_{C}$ that is closest to a root of $F$. We call $v_{C}$ the root of $C$ in $F$. We treat $H$ as a rooted forest, where each component $C$ of $H$ is rooted at $v_{C}$. For two vertices $u$ and $v$ in $F$, if $v$ is contained in the unique path from $u$ to a root of $F$, then we say that $u$ is a descendant of $v$ in $F$, and when additionally $u\neq v$ then $u$ is a strict descendant of $v$ in $F$. If $u$ is a strict descendant of $v$ and $uv\in E(F)$ then $u$ is a child of $v$. For a vertex $v$ in $F$ the subtree, $\color[rgb]{0.76,0.13,0.28}F(v)$, of $v$ is the subgraph of $F$ induced by all the descendants of $v$ in $F$ that is rooted at $v$. Let $X\subseteq V(F)$. We say that $x\in X$ is minimal in $X$ with respect to $F$ if no strict descendant of $x$ lies in $X$.
@@
For an indexed family $A:=(A_{i})_{i\in I}$ with index set $I$, we use $\color[rgb]{0.76,0.13,0.28}\pi_{i}(A)$ to denote the element $A_{i}$ of $A$ indexed by $i$. Let $I$ be a set and let $F_{i}$ be a forest for each $i\in I$. [^4] In our application, the forests in $(F_{i})_{i\in I}$ are pairwise vertex-disjoint, but this is not required in the following results. A sequence $(A_{i})_{i\in I}$ is a $(F_{i})_{i\in I}$-tuple if, for each $i\in I$, $A_{i}$ is a null graph or $A_{i}$ is a connected subgraph of $F_{i}$. An $(F_{i})_{i\in I}$-tuple $(A_{i})_{i\in I}$ is trivial if $A_{i}$ is a null graph for each $i\in I$ and is non-trivial otherwise. Let $A=(A_{i})_{i\in I}$ and $B=(B_{i})_{i\in I}$ be two $(F_{i})_{i\in I}$-tuples. We say that $A$ and $B$ are independent if $V(A_{i})\cap V(B_{i})=\emptyset$ for each $i\in I$. A set $\mathcal{A}$ of $(F_{i})_{i\in I}$-tuples is independent if the elements of $\mathcal{A}$ are pairwise independent.
@@
We make use of the following corollary, which is obtained immediately from Lemma˜14, below, by taking $m=k$ and $x_{1}=x_{2}=\cdots=x_{k}=1$.
@@
Let $c,k$ be positive integers. Let $F_{i}$ be a forest for each $i\in[c]$. For each $j\in[k]$, let $\mathcal{A}_{j}$ be a set of independent $(F_{i})_{i\in[c]}$-tuples with $|\mathcal{A}_{j}|\geq k^{c}$. Then, for each $j\in[k]$ there exists $B_{j}\in\mathcal{A}_{j}$ such that $B_{1},\ldots,B_{k}$ are pairwise independent.
To establish ˜13 we prove the following more general result, which is amenable to a proof by induction.
@@
Let $c,m,k$ be positive integers. Let $F_{i}$ be a forest for each $i\in[c]$. For each $j\in[m]$, let $\mathcal{A}_{j}$ be a set of independent $(F_{i})_{i\in[c]}$-tuples with $|\mathcal{A}_{j}|\geq m^{c-1}k$. Let $x_{1},\ldots,x_{m}$ be nonnegative integers with $x_{1}+\cdots+x_{m}=k$. Then, for each $j\in[m]$ there exist $A_{j}^{1},\ldots,A_{j}^{x_{j}}\in\mathcal{A}_{j}$ such that
@@
We call a tuple $(c,m,k,(F_{i})_{i\in[c]},(\mathcal{A}_{j})_{j\in[m]},(x_{j})_{j\in[m]})$ that satisfies the assumptions of the lemma a $(c,m,k)$-instance and we say that the sequence of sets in (10) satisfies this $(c,m,k)$-instance.
@@
Without loss of generality, we may assume that $x_{j}>0$ for each $j\in[m]$. To see why this is so, let $J:=\{j\in[m]:x_{j}>0\}$ and $m^{\prime}:=|J|$. Since $m^{\prime}\leq m$, $\mathcal{I}^{\prime}:=(c,m^{\prime},k,(F_{i})_{i\in[c]},(\mathcal{A}_{j})_{j\in J},(x_{j})_{j\in J})$ is a $(c,m^{\prime},k)$-instance with $x_{j}>0$ for each $j\in J$, and any sequence $(A_{j}^{\ell})_{j\in I,\,\ell\in[x_{j}]}$ that satisfies this instance also satisfies (10). From this point on we assume that $x_{j}>0$ for each $j\in[m]$.
@@
Without loss of generality, we assume that for all $j\in[m]$, $i\in[c]$, $A\in\mathcal{A}_{j}$, the tree $\pi_{i}(A)$ is non-null. Indeed, otherwise for all $j$, $i$, AA violating this rule, we introduce a new vertex in $F_{i}$ and make $\pi_{i}(A)$ be the graph that contains only this new vertex. Since this new vertex appears only in $\pi_{i}(A)$, the set $\mathcal{A}_{j}$ is still independent. Because $\mathcal{A}_{j}$ is independent for each $j\in[m]$, this assumption ensures that $\pi_{i}(A)\neq\pi_{i}(A^{\prime})$ for each $i\in[c]$, $j\in[m]$ and distinct $A,A^{\prime}\in\mathcal{A}_{j}$. This ensures that, for every non-empty $I\subseteq[c]$, every $j\in[m]$, and every $\tilde{\mathcal{A}}_{j}\subseteq\mathcal{A}_{j}$, we have $|\{(\pi_{i}(A))_{i\in I}:A\in\tilde{\mathcal{A}}_{j}\}|=|\tilde{\mathcal{A}}_{j}|$. In the following we will use this fact repeatedly without explicitly referring to it.
@@
The assumption above, that for all $j\in[m]$, $i\in[c]$, $A\in\mathcal{A}_{j}$, the tree $\pi_{i}(A)$ is non-null, has another consequence, which we explain now. Let $I$ be a non-empty subset of $[c]$, for each $j\in[m]$ let $\mathcal{B}_{j}$ be obtained from a subset of $\mathcal{A}_{j}$ by restricting the tuples to the coordinates in $I$. Let $k^{\prime}$ be a positive integer and let $x^{\prime}_{1},\dots,x^{\prime}_{m}$ be nonnegative integers with $x^{\prime}_{1}+\dots+x^{\prime}_{m}=k^{\prime}$. Let $\mathcal{I}_{\mathcal{B}}:=(|I|,m,k^{\prime},(F_{i})_{i\in I},(\mathcal{B}_{j})_{j\in[m]},(x^{\prime}_{j})_{j\in[m]})$ be a $(|I|,m,k^{\prime})$-instance. Then, for each $B\in\bigcup_{j\in[m]}\mathcal{B}_{j}$, the $(F_{i})_{i\in I}$-tuple $(\pi_{i}(B))_{i\in I}$ is non-trivial. Any two non-trivial independent $(F_{i})_{i\in I}$-tuples are necessarily distinct. It follows that any sequence that satisfies the instance $\mathcal{I}_{\mathcal{B}}$ is a sequence of $k^{\prime}$ distinct $(F_{i})_{i\in I}$-tuples. (Note that this would not necessarily be true if trivial tuples could appear.) Treating this sequence as a set does not change its size. Similarly, if there exist sets $\mathcal{B}_{1}^{\prime},\ldots,\mathcal{B}_{m}^{\prime}$ with $\mathcal{B}_{j}^{\prime}\subseteq\mathcal{B}_{j}$ and $|\mathcal{B}_{j}^{\prime}|=x_{j}^{\prime}$ for each $j\in[m]$ and such that $\bigcup_{j\in[m]}\mathcal{B}_{j}^{\prime}$ is independent and has size $k^{\prime}$ then there exists a sequence that satisfies the instance $\mathcal{I}_{\mathcal{B}}$. In this case, we say that $\mathcal{B}_{1}^{\prime},\ldots,\mathcal{B}_{m}^{\prime}$ satisfies $\mathcal{I}_{\mathcal{B}}$.
@@
With these two assumptions justified, we now proceed with the proof. Let $\mathcal{I}:=(c,m,k,(F_{i})_{i\in[c]},(\mathcal{A}_{j})_{j\in[m]},(x_{j})_{j\in[m]})$ be a $(c,m,k)$-instance satisfying the assumptions in the previous paragraphs. Our goal is to show the existence of a sequence of sets $\mathcal{A}^{\prime}_{1},\ldots,\mathcal{A}^{\prime}_{m}$ that satisfies the instance $\mathcal{I}$. We proceed by induction on $(c,k)$ in lexicographic order.
@@
Define $y_{j}:=x_{j}$ for each $j\in[m]\setminus\{\ell\}$ and $y_{\ell}:=x_{\ell}-1$. Now, (11) implies that $\mathcal{I}_{\mathcal{B}}:=(1,m,k-1,(F_{i})_{i\in[c]},(\mathcal{B}_{j})_{j\in[m]},(y_{j})_{j\in[m]})$ is a (1,m,k1)(1,m,k-1)-instance. By the inductive hypothesis, there exists $\mathcal{B}_{1}^{\prime},\ldots,\mathcal{B}_{m}^{\prime}$ that satisfy $\mathcal{I}_{\mathcal{B}}$. For each $j\in[m]\setminus\{\ell\}$, let $\mathcal{A}^{\prime}_{j}:=\mathcal{B}^{\prime}_{j}$ and let $\mathcal{A}^{\prime}_{\ell}:=\{\check{A}\}\cup\mathcal{B}^{\prime}_{\ell}$. Then the sequence $\mathcal{A}^{\prime}_{1},\ldots,\mathcal{A}^{\prime}_{m}$ satisfies $\mathcal{I}$. This completes the proof of the case $c=1$.
@@
is a $(1,m,m^{c-1}k)$-instance. Applying induction on $\mathcal{I}_{\mathcal{B}}$ we obtain a sequence of sets $\mathcal{B}_{1}^{\prime},\ldots,\mathcal{B}_{m}^{\prime}$ that satisfies the instance $\mathcal{I}_{\mathcal{B}}$.
For each $j\in[m]$, let $\mathcal{C}_{j}=\{(\pi_{2}(A),\ldots,\pi_{c}(A)):A\in\mathcal{A}_{j},\pi_{1}(A)\in\mathcal{B}^{\prime}_{j}\}$. Observe that $|\mathcal{C}_{j}|=|\mathcal{B}^{\prime}_{j}|=y_{j}=m^{c-2}k$. Therefore, the tuple $\mathcal{I}_{\mathcal{C}}:=(c-1,m,k,(F_{i})_{i\in\{2,\ldots,c\}},(\mathcal{C}_{j})_{j\in[m]},(x_{j})_{j\in[m]})$ is a (c1,m,k)(c-1,m,k)-instance. Applying induction yields $\mathcal{C}_{1}^{\prime},\ldots,\mathcal{C}_{m}^{\prime}$ that satisfies the instance $\mathcal{I}_{\mathcal{C}}$.
@@
There exists a function $\ell:\mathbb{N}^{2}\to\mathbb{N}$ such that, for all $k,c\in\mathbb{N}$ with $k,c\geq 1$, for every $(F_{i})_{i\in[c]}$ where $F_{i}$ is a forest for each $i\in[c]$, for every set $\mathcal{A}$ of non-trivial $(F_{i})_{i\in[c]}$-tuples either
@@
We will show that the lemma holds with
$\ell(k,c):=c\cdot\left(\prod_{j=1}^{c-1}2^{j!}\right)\,k^{c!}.$
In order to help the induction go through, we prove the lemma with the following slightly more precise property in place of ˜b:
  1. (b’)
    there exists a sequence of sets $(X_{i})_{i\in[c]}$ with $X_{i}\subseteq V(F_{i})$ and $|X_{i}|\leq\left(\prod_{j=1}^{c-1}2^{j!}\right)\,k^{c!}$ for each $i\in[c]$, and such that for each $(A_{1},\ldots,A_{c})\in\mathcal{A}$, there exists $i\in[c]$ with $X_{i}\cap V(A_{i})\neq\emptyset$.
Note that this implies that $\sum_{i\in[c]}|X_{i}|\leq c\cdot(\prod_{j=1}^{c-1}2^{j!})\,k^{c!}=\ell(k,c)$, so ˜b follows from ˜a.
@@
We may assume that $\pi_{i}(A)$ contains at least one vertex, for each $A\in\mathcal{A}$ and each $i\in[c]$. Indeed, otherwise we introduce a new vertex in $F_{i}$ and make $\pi_{i}(A)$ be the graph that contains only this new vertex. Since this vertex appears only in $\pi_{i}(A)$, this does not affect outcome ˜a. If this new vertex appears in the set $X_{i}$ promised by outcome ˜a, then we can remove it from $X_{i}$ and replace it with any other vertex in $\bigcup_{j\in[c]\setminus\{i\}}V(\pi_{j}(A))$. (The fact that $A$ is non-trivial for each $A\in\mathcal{A}$ ensures that $\bigcup_{j\in[c]\setminus\{i\}}V(\pi_{j}(A))$ contains at least one vertex.) This assumption ensures that for any non-empty $I\subseteq[c]$, and any $A\in\mathcal{A}$, $(\pi_{i}(A))_{i\in I}$ is non-trivial.
@@
Finally, we claim that we may assume that for each $i\in[c]$, $F_{i}$ is a binary tree, [^5] A binary tree is a rooted tree in which each vertex has at most two children. The root of a binary tree is not considered to be a leaf, even if it has no children., thanks to the following argument. For each $i\in[c]$, and each vertex $v$ in $F_{i}$ with children $u_{1},\ldots,u_{\beta}$, define $T_{i}(v)$ to be a binary tree rooted at $v$ and whose leaves are $u_{1},\ldots,u_{\beta}$, and let $T^{-}_{i}(v):=T_{i}(v)-\{u_{1},\ldots,u_{\beta}\}$. As expected, we choose these binary trees so that every two are vertex-disjoint except possibly for the one common vertex imposed by the definition; that is, given two distinct vertices $v,w$ in $F_{i}$, if $vw$ is not an edge in $F_{i}$ then $T_{i}(v)$ and $T_{i}(w)$ are vertex-disjoint, and if $v$ is the parent of $w$ then their only common vertex is $w$, which is then a leaf of $T_{i}(v)$. Now let $F_{i}^{\prime}:=\bigcup_{v\in V(F_{i})}T_{i}(v)$ and note that $F_{i}^{\prime}$ is a binary tree for each $i\in[c]$. For each $A\in\mathcal{A}$, define $A^{\prime}:=(A^{\prime}_{1},\ldots,A^{\prime}_{c})$ where $A^{\prime}_{i}:=F_{i}^{\prime}[\bigcup_{v\in V(\pi_{i}(A))}V(T^{-}_{i}(v))]$ for each $i\in[c]$ and define $\mathcal{A}^{\prime}:=\{A^{\prime}:A\in\mathcal{A}\}$. Then $\mathcal{A}^{\prime}$ is a set of non-trivial $(F_{i}^{\prime})_{i\in[c]}$-tuples. Observe that for all $A,B\in\mathcal{A}$, (1) AAand $B$ are independent if and only if $A^{\prime}$ and $B^{\prime}$ are independent; (2) if $A_{i}^{\prime}$ contains a vertex in $T_{i}^{-}(v)$ for some $i\in[c]$ and $v\in V(F_{i})$, then $A_{i}$ contains $v$. Now, if we know that Lemma˜15 holds in the special case where the forests are binary trees, we may proceed as follows: We apply the lemma to $(F^{\prime}_{i})_{i\in[c]}$ and $\mathcal{A}^{\prime}$. If this results in an independent set $I^{\prime}\subseteq\mathcal{A}^{\prime}$ of size $k$, then Item˜1 ensures that $I:=\{A\in\mathcal{A}:A^{\prime}\in I^{\prime}\}$ satisfies ˜a. If this results in sets $X_{1}^{\prime},\ldots,X_{c}^{\prime}$, then define $X_{i}:=\{v\in V(F_{i}):V(T_{i}^{-}(v))\cap X_{i}^{\prime}\neq\emptyset\}$ for each $i\in[c]$. Then Item˜2 ensures that the sets $X_{1},\ldots,X_{c}$ satisfy ˜a. Therefore, to prove Lemma˜15 in full, it is enough to prove it in the case of binary trees, as claimed. Assume thus from now on that $F_{i}$ is a binary tree for each $i\in[c]$.
We now proceed by induction on $c$. In order to avoid presenting an argument for $c=1$ that is a special case of the argument for $c>1$, we treat $c=0$ as the base case of the induction. In this case, the conclusion of the lemma is trivial since the empty sequence satisfies the conditions of outcome ˜a.
Now suppose that $c\geq 1$. Consider the following iterative process, that constructs a subset of vertices $\{v_{1},\ldots,v_{m}\}$ of $F_{1}$. Suppose that $v_{1},\ldots,v_{z-1}$ have already been defined, for some $z\geq 1$. If $r_{1}\in\{v_{1},\ldots,v_{z-1}\}$, then we set m:=z1m:=z-1 and stop the process. Otherwise, define